OpenStudy (anonymous):

Take the fraction and transform it into the sum or two fractions (method of partial fractions): (4x+3)/(x^2-x-6)

OpenStudy (cwrw238):

= 4x + 3 A B -------- = ---- + ----- (x - 3)(x + 2) (x - 3) (x + 2) multiply thru by (x-3)(x+2) 4x + 3 = A(x+2) + B(x-3) let x = 3 then 15 = 5A let x = -2 then -5 = -5B solve these 2 simple equations to get your partial fractions